Sorry, I had the wrong email address typo....
If Joe makes $3.00 more in Overtime then regular time, and if he were to work 5 hours straight time and 4 hours overtime for a total of $66.00. What is the straight time amount.
I figured out the answer is $6.00, but do not know how to write in equation form. Can you help with this?
Carolyn

Answer by edjones(8007)

let x be the straight time amt.
5x+4(x+3)=66
5x+4x+12=66
9x+12-12=66-12
9x=54
9x/9=54/9
x=$6.00
.
Check:
5*6+4*9
=30+36 true
.
Ed